OpenStudy (anonymous):

If the area of a rectangle 24a2b and the length is 8ab2, what would the width of the rectangle, given that width is found by dividing area by length? Simplify the answer.

OpenStudy (ranga):

Divide: 24(a^2)b / 8a(b^2) = 24/8 * a^2/a * b/b^2 = ?

OpenStudy (anonymous):

can you write out the equation?

OpenStudy (anonymous):

what does the thing mean thats inbetween the 24/8 and a^2/1?

OpenStudy (ranga):

\[\frac{ 24a^2b }{ 8ab^2 } = \frac{ 24 }{ 8 }*\frac{ a^2 }{ a}*\frac{ b }{ b^2 } = ?\]

OpenStudy (anonymous):

okay much better :)

OpenStudy (ranga):

The * means multiplication. sometimes using x is confusing because a lot of times x is a variable.

OpenStudy (anonymous):

thanks for that

OpenStudy (ranga):

yw. what do you get?

OpenStudy (anonymous):

one second...

OpenStudy (anonymous):

3a/b?

OpenStudy (anonymous):

Thank you for your help!!!

OpenStudy (ranga):

yes. you are welcome.

OpenStudy (anonymous):

that answer is wrong @ranga @zzkaitlyn

OpenStudy (anonymous):

Actually it was right